OpenStudy (darkigloo):

A radioactive element decays exponentially in proportion to its mass. One half of its original amount remains after 5570 years. If 10000 grams of the element are present initially, how much will be left after 1000 years?

OpenStudy (darkigloo):

I thought i had to do this but i'm wrong. The answer is supposed to be 8864g. y=y(o)e^kt k=-ln2/5570 y=10,000e^(k(1000)) y=10,000e^((-ln2/5570)(1000))